# OneDiff: A Generalist Model for Image Difference Captioning

### v1

@@ 5.4 Ablation Analysis
We first assess the efficacy of the proposed CST strategy and VDM module. We report the related results in Table 5. Specifically, on one hand, the CST strategy facilitates cost-effective and efficient adaptation for model to multi-image understanding domains, yielding improvements across all benchmarks. On the other hand, while the isolated implementation of the VDM module does not significantly outperform the baseline (which we attribute to insufficient training of the VDM due to the CST being disabled during the pretrain phase), the concurrent activation of both CST and VDM substantially boosts performance, significantly surpassing the performance of each of them when used independently. This demonstrates the synergistic benefits of the two proposed methods.
@@
In this section, we systematically analyze the impact of both realistic and synthetic data generated via our proposed pipeline, as indicated by the performance metrics outlined in the Table 6. Specifically, incorporating either type of data independently yield consistent performance gains across three benchmarks, which attests to the high quality and extensive distribution of our data. Moreover, utilizing both data types in tandem further enhance performance, underscoring the complementary nature of the two data patterns. Overall, the data we constructed have substantially augmented the existing IDC dataset, effectively mitigating the data scarcity issue in this field.
@@ 6 Conclusion
{ack}
By using the ack environment to insert your (optional) acknowledgements, you can ensure that the text is suppressed whenever you use the doubleblind option. In the final version, acknowledgements may be included on the extra page intended for references.

### v3

@@ 5.4 Ablation Analysis
We first assess the efficacy of the proposed CST strategy and VDM module. We report the related results in Tab. 2 with the 6-layer and 32-delta-token VDM. The hyper-parameter setting of VDM is discussed in our supplementary materials. Specifically, on one hand, the CST strategy facilitates cost-effective and efficient adaptation for model to multi-image understanding domains, yielding improvements across all benchmarks. On the other hand, while the isolated implementation of the VDM module does not significantly outperform the baseline (which we attribute to insufficient training of the VDM due to the CST being disabled during the pretrain phase), the concurrent activation of both CST and VDM substantially boosts performance demonstrating the synergistic benefits, significantly surpassing the performance of a single strategy. We opt for a more efficient language model in Tab. 3(a), reducing computational costs while maintaining commendable performance on Birds-to-Words. To further demonstrate the role of VDT, we randomly drop a portion of them during inference to test the impact on performance on Birds-to-Words in Tab. 3(b).
@@
In this section, we systematically analyze the impact of both realistic and synthetic data generated via our proposed pipeline, as indicated by the performance metrics outlined in Tab. 4. Specifically, incorporating either type of data independently yield consistent performance gains across three benchmarks, which attests to the high quality and extensive distribution of our data. Moreover, utilizing both data types in tandem further enhance performance, underscoring the complementary nature of the two data patterns. Overall, the data we constructed have substantially augmented the existing IDC dataset, effectively mitigating the data scarcity issue in this field.
@@ 6 Conclusion
{credits}
6.0.1 Acknowledgements
This research is supported by Artificial Intelligence-National Science and Technology Major Project (2023ZD0121200), the National Natural Science Foundation of China (No. 62436001, 62437001).
Appendix 0.A Appendix
In this appendix section, we provide following items:
  • Sec. 0.A.1 Ablation experiments onthe hyper-parameters settings of the VDM module in our OneDiff framework.
  • Sec. 0.A.2 The statistics of our proposed DiffCap Dataset, both real-image and synthetic-image samples from DiffCap and some potential failure cases.
0.A.1 Ablation
The hyper-parameter setting of VDM is explored in Tab. 5. For efficiency, we only train 3000 steps in Stage II when exploring the model design. The number of delta tokens varies in 16, 32, 64 and 128 with 6 or 12 Transformer layers. OneDiff is not quite sensitive to different hyper-parameters with relatively stable performances.
Table 5: Ablation of model design on Visual Delta Module (VDM). We conduct ablation analysis on three aforementioned benchmarks.
Layers Tokens Image-Editing-Request Spot-the-Diff Birds-to-Words
B C M R B C M R B C M R
6 16 28.1 98.8 24.0 51.0 9.8 43.1 13.6 32.4 25.0 19.6 26.0 46.4
6 32 28.4 99.7 23.4 51.4 10.1 47.6 13.8 32.4 25.9 20.6 25.9 46.8
6 64 28.7 100.9 24.0 52.0 10.5 45.5 14.3 33.1 24.4 18.3 25.1 46.2
6 128 26.0 93.1 22.7 49.7 10.1 45.7 13.5 32.4 24.7 19.5 25.9 45.9
12 16 22.9 94.7 23.0 50.8 9.6 42.8 13.8 32.1 26.1 21.8 25.9 46.3
12 32 23.4 99.9 23.5 50.7 9.2 40.1 13.2 31.6 24.9 22.2 25.7 46.6
12 64 27.4 97.6 23.6 50.3 9.8 44.5 13.6 32.5 23.9 16.6 25.8 45.0
12 128 24.0 101.6 23.9 52.1 9.7 44.4 13.9 32.4 24.8 20.6 26.0 46.9
0.A.2 Data
This section illustrates the statistics of DiffCap and the qualitative results of data construction.
0.A.2.1 Data Statistics.
Our DiffCap dataset consists of three components: collected datasets, generated data from real images and generated data from synthetic images. (See Tab. 6)
Table 6: Statistics of our DiffCap dataset. "Real" refers to our generated data from real images, and "Synthetic" refers to our generated data from synthetic images.
Component Sub-Component #Image-Pairs #Annotations
Collected Spot-the-Diff 11K 21K
CLEVR-Change 8K 44K
Image-Editing-Request 3.4K 4.2K
Birds-to-Words 14K 14K
NLVR2 86K 86K
Generated Real 54K 54K
Synthetic 139K 139K
Total 316K 363K
0.A.2.2 Data Samples.
More samples from DiffCap are displayed as qualitative results of date construction. (See Fig. 6 and Fig. 7)
Figure 6: More real-image samples of our DiffCap dataset.
Figure 7: More synthetic-image samples of our DiffCap dataset.
0.A.2.3 Potential Failure Cases
We present potential failure cases where the environment and objectives are roughly the same, yet multiple describable image differences exist, which could lead to confusion. (See Fig. 8)
Figure 8: Potential failure cases of our DiffCap dataset.